# Completely regular growth solutions to linear differential equations with exponential polynomials coefficients

### v1

@@ 4 Proofs
for $|z|\notin E\cup E_{\theta},z\in S_{j}(\varepsilon)$ as $z\rightarrow\infty$ in $\theta\leq\arg z\leq\theta+h,h>0$, and outside of a logarithmic semi-strip $0\leq\arg z-\theta<C\frac{\log^{+}|z|}{|z|^{1/p_{j}}}$. The last semi-strip occurs only if $\arg z=\theta$ is a Stokes ray in $S_{j}$, otherwise, we set $C=0$. Thus, it is natural to be inferred $f(z)$ is of completely regular growth on every ray $\theta\in S_{j}(\varepsilon)\backslash\{\arg z=\xi_{j1},\ldots,\arg z=\xi_{j_{m_{j}}}\},j=0,\ldots,N_{0},$ where $\{\arg z=\xi_{j1},\ldots,\arg z=\xi_{j_{m_{j}}}\}$ are the Stokes rays of equation (4.5) in $S_{j}(\varepsilon),m_{j}\in\mathbb{N}.$
Except for Stokes rays, we assert that $f(z)$ is of completely regular growth on every ray $\theta=\arg z$ except for the critical rays $\arg z=\eta_{j},j=0,\ldots,N_{0}.$ Since
$S_{j}(\varepsilon)=\{z|\eta_{j-1}+\varepsilon\leq\arg z\leq\eta_{j}-\varepsilon\},$
if there exists a $\eta_{j}-\varepsilon<\theta_{0}<\eta_{j}$ such that $f(z)$ is not of completely regular growth on the ray $\arg z=\theta_{0},$ we can always find a $\varepsilon_{1}<\eta_{j}-\theta_{0},$ with $f$ of completely regular growth on the ray in
$S_{j}(\varepsilon_{1})=\{z|\eta_{j-1}+\varepsilon_{1}\leq\arg z\leq\eta_{j}-\varepsilon_{1}\},$
a contradiction.
That means the set of arguments of completely regular growth
$\mathbb{M}=(0,2\pi]\backslash\{\eta_{0},\ldots,\eta_{N_{0}},\xi_{01},\ldots,\xi_{0{m_{0}}},\ldots,\xi_{N_{0},1},\ldots,\xi_{N_{0}{{m_{N_{0}}}}}\}.$
@@
Then, by Lemma 2.1, $f(z)$ is of completely regular growth on the set of rays with arguments $\overline{\mathbb{M}}=(0,2\pi]$. $\Box$
@@
Furthermore, assuming there is a $A\in\mathbb{R}$ such that for a sufficiently large sequence $|z_{\mu}|=r_{\mu}$,
@@
    for a $B\in\mathbb{R}$, except for a set $E$ in $r=|z|$ of finite linear measure, since $f^{(m)}/f$ is of polynomial growth except for some set $E$ in $r=|z|$ of finite linear measure, see Gundersen [4, Corollary 3], such that
@@
    Now, $\alpha$ can be chosen from the interval $(1-1/s,1)$ and
@@
    If $\overline{q}_{t}$ lies on $s_{p+1},t\neq p.$ As in Case 2, we have $\kappa_{t}(z_{\mu})=o(1).$
@@
$\left|\sum_{j=1}^{k}H_{j}(z_{\mu})e^{q_{j}z_{m}u^{s}}\right|\geq c(1+o(1))e^{Ar_{\mu}^{s-1}-|q_{p}z_{\mu}^{s}|},\enspace|z_{\mu}|\notin E,$
@@
By the similar proof of theorem 3.2 and Lemma 2.4, we afford
@@
5 A blick of the boundary of function class c.r.g.
Observing the two main Theorems 3.2 and 3.6, the solutions of completely regular growth have asymptotic expressions

### v2

@@ 4 Proofs of main results
for $|z|\notin E_{\theta},z\in S_{j}(\varepsilon)$ as $z\rightarrow\infty$ in $\theta\leq\arg z\leq\theta+h,h>0$, and outside of a logarithmic semi-strip $0\leq\arg z-\theta<C\frac{\log^{+}|z|}{|z|^{1/p_{j}}}$. The last semi-strip occurs only if $\arg z=\theta$ is a Stokes ray in $S_{j}$, otherwise, we set $C=0$. Thus, it is natural to be inferred $f(z)$ is of completely regular growth on every ray $\theta\in S_{j}(\varepsilon)\backslash\{\arg z=\xi_{j1},\ldots,\arg z=\xi_{j_{m_{j}}}\},j=0,\ldots,N_{0},$ where $\{\arg z=\xi_{j1},\ldots,\arg z=\xi_{j_{m_{j}}}\}$ are the Stokes rays of equation (4.5) in $S_{j}(\varepsilon),m_{j}\in\mathbb{N}.$
If $f$ has infinitely many zeros in $S_{j}(\varepsilon)$, the associated exceptional set $\tilde{B}$ consists of countably many closed disks constructed via Cartan’s lemma. Following the notation in (4.8), the complex plane is partitioned into annular strips $\alpha^{\nu}\leq|z|\leq\alpha^{\nu+1}$ with $\alpha>1$ and $\nu\geq n_{0}$. Within each strip, a finite collection of disks $B_{\nu,t}=B(\zeta_{\nu,t},r_{\nu,t})$, $t=1,\dots,p_{\nu}$, covers the zeros of $f$ in that strip. The total linear measure of these disks is finite, i.e., $\sum_{\nu,t}4\alpha^{-\varepsilon_{0}\nu}<\infty$. The classical asymptotic theorems of Wasow and Steinmetz require holomorphy on a full sector, which is not guaranteed on $S_{j}^{*}$ due to the presence of these disks.
To overcome this, we employ the modified Wasow theorem (Theorem 6.5) on a subregion $S_{0}\subset S_{j}^{*}\subset S_{j}(\varepsilon)$ that satisfies the uniform interior condition. The construction of $S_{0}$ involves slightly enlarging each exceptional disk: define $B^{0}_{\nu,t}=B\!\left(\zeta_{\nu,t},r_{\nu,t}+\frac{\alpha_{0}}{|\zeta_{\nu,t}|}\right)$ for a fixed $\alpha_{0}>0$, and set $S_{0}=S_{j}(\varepsilon)\setminus\bigcup_{\nu,t}B^{0}_{\nu,t}$.
The crucial observation is that the total linear measure of the enlarged disks remains finite. Indeed, since the original exceptional set has finite linear measure, we have $\sum_{\nu}4\alpha^{-\varepsilon_{0}\nu}<\infty$. Moreover, because of $\alpha^{\nu}\leq|\zeta_{\nu,t}|\leq\alpha^{\nu+1}$, by [13, eq. (7.2),(7.6),(7.7)], we denote $p_{\nu}$ as the number of cartan disks which intersects with the annulus, and $n(r,1/f)$ as the counting function of zeros of $f$ inside the disk $B(0,r)$, and then we get
$p_{\nu}\leq n\left(\alpha^{\nu+2},\frac{1}{f}\right)\leq\frac{12}{\log\alpha}T(\alpha\alpha^{\nu+2},f)<\frac{12}{\log\alpha}\alpha^{(\nu+3)(\rho+\varepsilon_{0})},$
so the sum of diameter-difference between $B_{\nu,t}$ and $B_{\nu,t}^{\prime}$ is
$\sum_{\nu}\sum_{t=1}^{p_{\nu}}\left(\frac{2\alpha_{0}}{|\zeta_{\nu,t}|}\right)<\sum_{\nu}\sum_{t=1}^{p_{\nu}}\left(2\alpha_{0}\alpha^{-\nu}\right)<C\sum_{\nu}\alpha^{\nu(\rho+\varepsilon_{0}-1)},~C>0.$ (4.11)
Consequently, when $\rho(f)<1,$
$\sum_{\nu}\left(4\alpha^{-\varepsilon_{0}\nu}+\frac{\alpha_{0}}{|\zeta_{\nu,t}|}\right)<\infty,$
so the enlarged exceptional set $\bigcup_{\nu,t}B^{\prime}_{\nu,t}$ also has finite linear measure. Therefore, the subregion $S_{0}$ has full measure in $S_{j}(\varepsilon)$ and satisfies the uniform interior condition: for every $z\in S_{0}$, the disk $B(z,\alpha/|z|)$ is contained in $S_{j}(\varepsilon)\setminus\bigcup_{\nu,t}B_{\nu,t}=S^{*}$.
Since the radius $\alpha/|z|$ of interior condition is sharp in [38, Theorem 8.8] by the nature of Cauchy integral formula, we can not reduce it to $O(1/|z|^{N})$ for some integer $N$, to allow a uniform termwise-differentiable asymptotic form of $f^{\prime}(z)$ on such $S_{0}$ outside “much smaller” disks, in Corollary 6.3.
However, when $\rho(f)\geq 1$ and $n(r,1/f)$ grows extremely fast as r growing, the sum (4.11) may not be coved by a finite linear measure. In this case, we can deal with the exceptional disks on the $\xi$-plane under the tranformation $\xi=z^{n_{0}}$, $n_{0}=\lfloor\rho+1\rfloor$.
Let $g(\xi)=f(z)=f(\xi^{1/n_{0}})$. Without generality, we fix $j=1$ and abberviate the above symbols with the subscript $j$. We choose a single-valued analytic branch of $\xi^{1/n_{0}}$ corresponding to the sector $S(\varepsilon)$, and denote by $S^{\prime}(\varepsilon)$ the image of $S(\varepsilon)$ under this mapping. The exceptional set $\tilde{B}$ is mapped to $\tilde{B}^{\prime}$,
$\tilde{B}^{\prime}:=\bigcup_{\nu=n_{0}}^{\infty}\bigcup_{t=1}^{p_{\nu}}B_{\nu,t}^{\prime}=\bigcup_{\nu=n_{0}}^{\infty}\bigcup_{t=1}^{p_{\nu}}B(\xi_{\nu,t},r_{\nu,t}^{\prime}),~n_{0}>N,$
and we set ${S^{*}}^{\prime}=S^{\prime}(\varepsilon)\setminus\tilde{B}^{\prime}$. Substituting $z=\xi^{1/n_{0}}$ into the asymptotic differential equation (4.9) and using the chain rule, we obtain a new linear differential equation for $g(\xi)$:
$g^{(n)}(\xi)+\sum_{k=1}^{n-1}Q_{k}(\xi)g^{(k)}(\xi)+\Bigl(Q_{0}(\xi)+o\bigl(\xi^{-(N_{1}+\deg\{a_{n}\})/n_{0}}\bigr)\Bigr)g(\xi)=0,$ (4.12)
where the coefficients $Q_{k}(\xi)$ are holomorphic in $S^{*}$$^{\prime}$ and admit asymptotic expansions of the form
$Q_{k}(\xi)\sim\xi^{\delta_{k}}\sum_{\nu=0}^{\infty}\frac{c_{k,\nu}}{\xi^{\nu/n_{0}}},\qquad\xi\to\infty,\;\xi\in{S^{*}}^{\prime},$ (4.13)
with rational exponents $\delta_{k}$ and constants $c_{k,\nu}\in\mathbb{C}$. In parallelled, we can construct $S_{0}^{\prime}$ involves slightly enlarging each exceptional disk: define ${B^{*}_{\nu,t}}=B\!\left(\xi_{\nu,t},r_{\nu,t}^{\prime}+\frac{\alpha_{0}}{|\xi_{\nu,t}|}\right)$ for a fixed $\alpha_{0}>0$, and set $S_{0}^{\prime}=S^{\prime}(\varepsilon)\setminus\bigcup_{\nu,t}B^{*}_{\nu,t}$. It is easy to check the uniform interior condition: there exists a constant $\alpha^{\prime}>0$ such that for every $\xi\in S_{0}^{\prime}$ the disk $B(\xi,\alpha^{\prime}/|\xi|)$ is contained in $S^{*}$$^{\prime}$. Therefore, the modified version of Wasow’s asymtotic theorem (Theorem 6.5) can be similarly used on the $\xi$-plane.
We will not repeat the process to get a similar estimate of $\log|g(\xi)|$ outside some exceptional set via Steinmetz’s method, compared to (4.10). Instead, our main objective is to demonstrate that through transformation $\xi=z^{n_{0}}$, the expansion of the diameters of these disks is measurable. Since the pre-image of each $B^{*}_{\nu,t}$} is $B\!\left(\zeta_{\nu,t},r_{\nu,t}+\frac{\alpha_{1}}{|\zeta_{\nu,t}|^{n_{0}}}\right)\supset B_{\nu,t}$, the sum of diameter-difference in (4.11) becomes
$\sum_{\nu}\sum_{t=1}^{p_{\nu}}\left(\frac{2\alpha_{1}}{|\zeta_{\nu,t}|}\right)<\sum_{\nu}\sum_{t=1}^{p_{\nu}}\left(2\alpha_{1}\alpha^{-n_{0}\nu}\right)<C\sum_{\nu}\alpha^{\nu(\rho-n_{0}+\varepsilon_{0})},~C>0.$
In conclusion, no matter how large is $\rho(f)$, we can treat the cartan disks in Gundersen’s estimate and following operations as a finite linear measure.
Thus, we may apply Theorem 6.5 to the linear differential equation on $S^{*}$, obtaining the asymptotic representation for $f$ as in (4.10) for $z\in S_{0}$ (hence for all $z\in S_{j}(\varepsilon)$ outside a set of zero relative linear measure). On such punctured sector $\B~S_{j}(\varepsilon)\backslash\tilde${B}, Steinmetz’s analysis framework (proof of Lemma 2.2) can still be generalized. We can then only replace the exceptional condition $|z|\notin E_{\theta}$ with
$|z|\notin E_{\theta}~\text{as well as}~z\notin\tilde{B}$,
and maintain that (4.10) holds. Since the definition of c.r.g. with exceptional set $C_{0}$-set and $E_{0}$-set are essentially equivalent, referring to Section 2.1, here we simplify the discussion about them. $\Box$
4.2 Proof of Theorem 3.5
Proof. To deal with the case $z\in T_{j}(\varepsilon)$, without generality, we set $j=1$. Due to Corollary 3.3, the normal form (3.3) possesses asymtptotic form (3.5) in $T_{1}(\varepsilon)$. Dividing $f$ by both sides of (3.5), we get
$\sum_{\overline{\lambda}_{k}\in s_{1}}e^{\lambda_{k}z}\sum_{m=0}^{n}c_{m,k}z^{d_{m,k}}(1+\varepsilon(z))\frac{f^{(m)}}{f}=0.$ (4.14)
First, we consider $z\in T_{1}(\varepsilon)\cap U_{1}$. Contrasting Lemma 3.1 with Corollary 3.4, the summation of (4.14) can be divided by
$\begin{split}&e^{\omega_{1}z}\sum_{m=0}^{n}c_{m,1}z^{d_{m,1}}(1+\varepsilon(z))\frac{f^{(m)}}{f}+\sum_{\tau_{p,m}\notin\tilde{s}_{1}}e^{\lambda_{p}z}\sum_{m=0}^{n}c_{m,p}z^{d_{m,p}}(1+\varepsilon(z))\frac{f^{(m)}}{f}\\ +&\sum_{\tau_{p,m}\in\tilde{s}_{1}}e^{\lambda_{p}z}\sum_{m=0}^{n}c_{m,p}z^{d_{m,p}}(1+\varepsilon(z))\frac{f^{(m)}}{f}=0.\end{split}$ (4.15)
We denote
$L_{k}^{*}(f)=\sum\limits_{m=0}^{n}c_{m,k}z^{d_{m,k}}(1+\varepsilon(z))\dfrac{f^{(m)}}{f},$
so (4.15) is abbreviated to
$e^{\omega_{j}z}L_{1}^{*}(f)+\sum_{\tau_{p,m}\notin\tilde{s}_{1}}e^{\lambda_{p}z}L_{p}^{*}(f)+\sum_{\tau_{p,m}\in\tilde{s}_{1}}e^{\lambda_{p}z}L_{p}^{*}(f)=0.$
Set $\max_{m}\{d_{m,1}\}$ as $d_{M}$. Then, we claim that the equality below is true:
$\frac{L_{1}^{*}(f)}{z^{d_{M}}}=o\left(z^{N_{1}}\right),\kern 5.0pt\text{for some}\kern 5.0ptN_{1}\in\mathbb{N},\kern 5.0ptz\in T_{1}(\varepsilon)\cap U_{1},\kern 5.0pt|z|\notin E,$ (4.16)
where $E$ is a set of finite linear measure.
Otherwise, there is a sequence $\{z_{i}\},z_{i}\in T_{1}(\varepsilon)\cap U_{1},|z_{i}|\notin E$, with $\lim\limits_{i\rightarrow\infty}z_{i}=\infty,$ such that
$\left|\sum\limits_{m=0}^{n}c_{m,1}z_{i}^{d_{m,1}-d_{M}}(1+\varepsilon(z_{i}))\frac{f^{(m)}(z_{i})}{f(z_{i})}\right|>{C}{|z_{i}|}^{N_{1}},$ (4.17)
for any $C,N_{1}\in\mathbb{N}$. [13, Corollary 3] point out that if we set $\rho=\rho(f)$ as growth order of $f$, there exists $\varepsilon_{0}>0$, such that
$\left|\frac{f^{(k)}}{f}\right|<{|z|^{k(\rho+\varepsilon_{0})}},\quad z\notin E.$ (4.18)
It holds for $\tau_{p,m}\in\tilde{s}_{1}$ that
$\kappa_{p}(z_{i}):=\left|\frac{e^{\lambda_{p}z_{i}}\left(\sum\limits_{m=0}^{n}c_{m,p}z_{i}^{d_{m,p}}(1+\varepsilon(z_{i}))\dfrac{f^{(m)}}{f}\right)}{z^{d_{M}}e^{\omega_{1}z_{i}}\left(\sum\limits_{m=0}^{n}c_{m,1}z_{i}^{d_{m,1}-d_{M}}(1+\varepsilon(z_{i}))\dfrac{f^{(m)}}{f}\right)}\right|=o(1),~|z_{j}|\notin E,$
for the reason that $z^{d_{m,p}-d_{M}}e^{\Re{(\lambda_{k}-\omega_{1})}z}$ decreases exponentially by Corollary 3.4, and the remained term in the absolute value is up to a polynomial rate of growth by (4.17) and (4.18). For $\tau_{p,m}\notin\tilde{s}_{1}$, by Corollary 3.4 and [13, Corollary 3],
$\kappa_{p}(z_{i})=o\left(|z_{i}|^{-v+n(\rho+\varepsilon_{0})-N_{1}}\right)=o(1),$
because of the existence of $N_{1}>n(\rho+\varepsilon_{0})-v$. Then for both of $\tau_{p,m}\in\tilde{s}_{1}$ and $\tau_{p,m}\notin\tilde{s}_{1}$,
$e^{\lambda_{p}z_{i}}\left(\sum\limits_{m=0}^{n}c_{m,p}z_{i}^{d_{m,p}}(1+\varepsilon(z_{i}))\dfrac{f^{(m)}}{f}\right)=o\left(e^{\omega_{1}z_{i}}\left(\sum\limits_{m=0}^{n}c_{m,1}z_{i}^{d_{m,1}}(1+\varepsilon(z_{i}))\dfrac{f^{(m)}}{f}\right)\right),$
which leads that for large $|z_{i}|\notin E$,
$z_{i}^{d_{M}}e^{\omega_{1}z_{i}}\left(\sum\limits_{m=0}^{n}c_{m,1}z_{i}^{d_{m,1}-d_{M}}(1+\varepsilon(z_{i}))\dfrac{f^{(m)}}{f}\right)(1+o(1))=0,$
a contradiction to the assumption (4.17).
Therefore, for $z\in T_{1}(\varepsilon)\cap U_{1},|z|\notin E$, (4.16) can be transformed into an equation
$f^{(n)}+P_{n-1}f^{(n-1)}+\cdots+P_{1}f^{\prime}+\left(P_{0}+o\left(z^{N_{1}-d_{n,1}}\right)\right)f=0,$ (4.19)
where $P_{m}=\frac{c_{m,1}z^{d_{m,1}}}{c_{n,1}z^{d_{n,1}}}(1+\varepsilon(z)),m=0,\ldots,n-1$. It is deduced that for $z\in T_{1}(\varepsilon)\cap U_{1},$ $|z|\notin E$, every $P_{m}$ can be asymptotically expanded as
$P_{m}(z)\sim z^{d_{m,1}-d_{n,1}}\sum_{v=0}^{\infty}\frac{b_{m,v}}{z^{v}},\quad|z|>|z_{0}|~\text{large enough},$ (4.20)
for $b_{m,v}\in\mathbb{C},m=1,\ldots,n-1.$ Taking a similar analysis to the proof section 4.1, using the modified version of Lemma 2.4 on a toplogical hull, Theorem 6.5, there exists a polynomial $G_{\theta}$ in $z^{1/p_{1}}$, $p_{1}\in\mathbb{N},$ a set of zero relative measure $E_{\theta}$ such that for $|z|\notin E_{\theta}$,
$z\in\{z|\theta\leq\arg z\leq\theta+h,0<h<\varepsilon_{1}\varepsilon,\varepsilon_{1}>0\}\subset T_{1}(\varepsilon)\cap U_{1}~\text{as}~|z|\rightarrow\infty,$
and outside of a logarithmic semi-strip $0\leq\arg z-\theta<C\frac{\log^{+}|z|}{|z|^{1/p}}$, it holds that
$\log|f(z)|=\Re G_{\theta}\left(z^{\frac{1}{p_{1}}}\right)+O(\log|z|).$ (4.21)
For the case $z\in T_{1}(\varepsilon)\cap U_{K^{\prime}}$, we can take parallel consideration to the case $z\in T_{1}(\varepsilon)\cap U_{1}$. since $j$ was arbitarily chosen as 1 in the above paragraph, we have an estimate of $\log|f|$ in the form of (4.21) for every $j=0,1\ldots,N_{0}$. To denote the general symbols, we write $U_{1}$ and $U_{K^{\prime}}$, the case to $j=1$, as $U_{j,1}$ and $U_{j,K^{\prime}_{j}}$, the case to arbitrary $j$. The area
$T_{j}(\varepsilon)\cap U_{j,1}\bigcup S_{j}(\varepsilon)\bigcup T_{j-1}(\varepsilon)\cap U_{j-1,K^{\prime}_{j-1}},$
which is a continuation of the sector $S_{j}(\varepsilon)$, contains every ray in $\overline{S}_{j}$ (see the definition in Section 3.1) originating from 0 when $|z|\to\infty$, besides the two critical rays $\arg z=\eta_{j-1}$ and $\arg z=\eta_{j}$.
In conclusion, combining (4.10) and (4.21), for $|z|\notin$ a $E_{0}$-set (or $z\notin$ a $C_{0}$-set) and $|z|\to\infty$, we get an estimate of
$\log|f|=\Re G_{\theta}\left(z^{\frac{1}{p_{\theta}}}\right)+O(\log|z|)$
in every narrow sector $\{z|\theta\leq\arg z\leq\theta+h,h>0\}$ outside some possibly existing logarithmic semi-strips $0\leq\arg z-\theta<C\frac{\log^{+}|z|}{|z|^{1/p_{\theta}}},$ when $\arg z=\theta$ is a Stokes ray, and outside every critical ray $\arg z=\eta_{j}$. That means the set of arguments of c.r.g. is
[…]
@@
where $\xi_{j,k}$, $k=1,\ldots,m_{0},$ are the Stokes rays insides $\overline{S}_{j}$. Then, by Lemma 2.1, $f(z)$ is of c.r.g. on the set of rays with arguments $\overline{\mathbb{M}}=(0,2\pi]$. $\Box$
@@
Furthermore, we claim there is a $A\in\mathbb{R}$ such that for a sufficiently large sequence $|z_{\mu}|=r_{\mu}$,
@@
    for a $B\in\mathbb{R}$, except for a set $E$ in $r=|z|$ of finite linear measure, since $f^{(m)}/f$ is of polynomial growth except for some set $E$ in $r=|z|$ of finite linear measure, see Gundersen [13, Corollary 3], such that
@@
    Now, $\alpha$ can be chosen from the interval (11/s,1)(1-1/s,1) and
@@
    If $\overline{q}_{t}$ lies on $L_{p+1},t\neq p.$ As in Case 2, we have $\kappa_{t}(z_{\mu})=o(1).$
@@
$\left|\sum_{j=1}^{k}H_{j}(z_{\mu})e^{q_{j}z_{m}u^{s}}\right|\geq c(1+o(1))e^{Ar_{\mu}^{s-1}-|q_{p}z_{\mu}^{s}|},\kern 5.0pt|z_{\mu}|\notin E,$
@@
By the similar proof of theorem 3.2 and Lemma 2.4 (when Lemma 2.4 lose effect for infinite many zeros of $f$ in each sector, we take the same discussion with (4.10) into account, which is abbreviated in this part), we afford
@@
In this proof, we will not repeat the analysis of the Cartan discs for the uniform interior expansion. It is always contained within a finite linear measure. See the ending of section 4.1.
@@
5 A dynamical aspect of c.r.g. solutions of LDEs
5.1 A nested concept – exponential polynomial type
Observing the two main Theorems 3.2 and 3.9, the solutions of completely regular growth have asymptotic expressions